There are 9 counters in a bag. 7 of the counters are green. 2 of the counters are blue. Ria takes at random two counters from the bag. Work out the probability that Ria takes one counter of each colour. You must show your working.

Probability that Ria takes Green then Blue is 7/9 * 2/8 = 14/72. Probability that Ria takes Blue then Green is 2/9 * 7/8 = 14/72. Added together we get 28/72 = 7/18.

ZG

Related Maths GCSE answers

All answers ▸

y=mx+c. If m=5, x= -3 and c=1, find y


Rearrange the equation y=3x+2 to make x the subject.


Work out the value of 2^4


The equation of line L1 is y=3x-5. The equation of line L2 is 2y-6x+5=0. Show that these two lines are parallel.